OpenStudy (imnotgoodatmath):

Janelle bought a total of 6 sodas and 6 hotdogs for herself and five friends at the ballpark. She spent a total of $23.70. Part 1: Write the equation for this word problem. Be sure to identify your variables. Part 2: If each hot dog cost $2.25, solve the equation to find the cost of each soda. Show your work.

OpenStudy (crashonce):

let x = soda and y=hotdogs 6x+6y=$23.70

OpenStudy (crashonce):

Because y=$2.25 therefore 6x=23.70-6x2.25 therefore 6x=1.70